I have been working and reworking this problem all night and day; please help if possible! Thanks in advance
Solve by any method and show equations and major steps:
Computerwarehouse.com charges $3 for shipping orders up to 10 lb, $5 for orders from 10 lb to 15 lb, and $7.50 for orders of 15 lb or more. One day shipping charges for 150 orders totaled $680. The number of orders under 10 lb was three times the number of orders weighing 15 lb or more. Find the number of packages shipped at each rate.

Answer by Paul(988)

Under 10lb = 3x
15 lb = x
from 10lb 10 15lb = y
3x+x+y=150
y=150-4x
Max value is 150:
Under 10lb the charge is $3
From 10lb to 15lb the charge is $5
More that 15lb the charge is $7.5
Total profit: $680
Equation:
3(3x)+5(y)+7.5(x)=680
9x+5y+7.5x=680
16.5x+5(150-4x)=680
16.5x-20x=680-750
-3.5x=-70
x=20
3(20)=60
y=150-4(20)
y=70
Hence, for 15lb or more their were 20 pakages, for 10lb to 15lb their were 70 pakages and for under 10lb their were 70 pakages.
Paul.